Write each algebraic fraction in its lowest terms: (2a³ - 4a² - 16a) ÷ (2a³ - 6a² - 20a)
Answer by It is costly(175)



(2a³-4a²-16a)/(2a³-6a²-20a)=
2a(a²-2a-8)/2a(a²-3a-10)=

(a+2)(a-4)/(a-5)(a+2)=
(a-4)/(a-5)